A new cruise ship line has just launched 3 new ships: the Pacific Paradise, the Caribbean Paradise, and the Mediterranean Paradise. The Caribbean Paradise has 33 more deluxe staterooms that the Pacific Paradise. The Mediterranean Paradise has 36 fewer deluxe staterooms that 4 times the number of deluxe staterooms on the Pacific Paradise. Find the number of deluxe staterooms for each of the ships if the total numbers of deluxe staterooms for the three ships is 837.
1
Expert's answer
2012-09-25T11:11:36-0400
Let's make such denotations:
Pacific Paradise - P rooms
Caribbean Paradise - C rooms
Mediterranean Paradise - M rooms

Let's formalize the problem statements now:

The caribbean paradise has 33 more deluxe staterooms than the Pacific Paradise, so

C = P + 33.

The Mediterranean Paradise has 36 fewer deluxe staterooms than four times the number of deluxe staterooms of the pacific paradise, so

M = 4P - 36.

At last, the total number of deluxe staterooms for the three ships is 837, so

P + C + M = 837.

Here we got the system of equations:

C = P + 33,& (1)
M = 4P - 36,& (2)
P + C + M = 837. (3)

Let's solve it.
substituting C from (1) and M from (2) to (3) we obtain:
P + P + 33 + 4P - 36 = 837 ==> 6P = 840 ==> P = 140.

Then,
C = P + 33 = 140 + 32 = 172
and
M = 4P - 36 = 4*140 - 36 = 524.

So, Pacific Paradise has 140 rooms, Caribbean Paradise 172 rooms and Mediterranean Paradise 524 rooms.